OpenStudy (help!!!!):

Help me simplify this @study100 @mathmate @jim_thompson5910

OpenStudy (help!!!!):

\[\frac{6\sqrt{10}}{4-\sqrt{2}}\]

OpenStudy (help!!!!):

Help me out with the steps

OpenStudy (help!!!!):

@thomaster

OpenStudy (anonymous):

first, I think that you would want to get the radical out of the denominator. Start with multiplying top and bottom by the conjugate of the denominator.

OpenStudy (anonymous):

doing that you'd get (24sqrt10+sqrt2*squrt360)/14

OpenStudy (anonymous):

this is more concise an accurate. http://www.wolframalpha.com/input/?i=%286%E2%88%9A10%29%2F%284-%E2%88%9A2%29

OpenStudy (anonymous):

can you please take a stab at the question I posted? thanks.

OpenStudy (midhun.madhu1987):